5. A balloon has a volume of 2.25 L when the temperature is 40 degree celsius. What will be the volume if the balloon is cooled to 12 degree celsius? Assume the pressure to be constant.


1
Expert's answer
2022-05-17T04:27:04-0400

Given:

V1 = 2.25 L

T1 = 40°C + 273.15 = 313.15 K

P1 = P2 = constant

V2 = unknown

T2 = 12°C + 273.15 = 285.15 K


Formula: V1/T1 = V2/T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the resulting volume, solve the equation for V2:

V2 = V1T2 / T1

V2 = (2.25 L × 285.15 K) / (313.15 K) = 2.05 L

V2 = 2.05 L


Answer: 2.05 L